OpenStudy (benlindquist):

2.14 + 12.18 = 0. how to do the steps?

pooja195 (pooja195):

Add 2.14+12.18

OpenStudy (benlindquist):

when i looked for help on another site, they said not to and change it around

pooja195 (pooja195):

Well you add those but you end up with a false statement.... Is there supposed to be a variable?

OpenStudy (benlindquist):

2.14x - 12.18 = -5.76,

pooja195 (pooja195):

What do you think the first step is?

OpenStudy (benlindquist):

you change them around, correct?

pooja195 (pooja195):

Yes

pooja195 (pooja195):

2.14x - 12.18+12.18 = -5.76+12.18

OpenStudy (benlindquist):

2.14x - 24.36 = 6.42

pooja195 (pooja195):

- + + gets cancelled put 2.14x=6.24

OpenStudy (benlindquist):

then divide

pooja195 (pooja195):

Yes

pooja195 (pooja195):

x=?

OpenStudy (benlindquist):

3.4

pooja195 (pooja195):

- + + gets cancelled put 2.14x=6.24 I meant 6.42 6.42/2.14=?

OpenStudy (benlindquist):

x = 3

pooja195 (pooja195):

Well done =)